A football is kicked with a speed of 18 m/sec an a
Answer #1

7) a) horizontal component of the initial velocity is given by

V = UCos\Theta--------(1)

here from eqn (1) we have

U = V/Cos\Theta

   = 18/Cos(65)

   = 18/0.4226----(2)

U = 42.59m/s

vertical component of the initial velocity is

U' = V / Sin\Theta

   =18/Sin(65)

=18/0.9060

=19.86m/s----------(3)

b) average velocity along the horizontal and vertical is given by

U1 = (U + U') / 2-------(4)

      = (42.59 + 19.86)/2

U1 = 31.22m/s------(5)

V = U1 -gt-------(6)

18 = 31.22 - (9.8)(t)

t = 13.22 / 9.8 = 1.34sec

football takes 1.34sec in air

c) V^2 = U^2 + 2gS

      (18)^2     = (31.22)^2 + 2(9.8)S

S = 33.19cm

8) the ball falls vertically through a distance

S = 1-0.8 = 0.2m

let 'V' be the velocity at the time of falling WHEN U =0

V^2 = U^2 + 2gS

V^2 = 0 + 2(9.8)(0.2)

V = 1.97m/s

velocity V = S/t

3 = 0.2/t

t = 0.06sec

acceleration a = (V - U)/t

a = (3-1.97)/0.06

= 17.16m/s^2

force   F = mxa

             = (0.5)(17.16)

    F = 8.58N

IMPULSE= Ft

           = (8.58)(0.06)

        = 0.514Ns